Gas systems project

Europe/Zurich
155/R-047 (CERN)

155/R-047

CERN

10
Show room on map
Beatrice Mandelli (CERN)

GAS M&O meeting 30 April 2020

Present: Beatrice, Roberto, Andrea, Patrick, Kacper, LPh, Fred, Michal, Mara, Gianluca, Sabino, Fabio, Federico, Leonardo

Indico: https://indico.cern.ch/event/912569/

 

 

CHECK Gas Systems

 

ALICE #2423

TOF

  • GC analysis: SF6 at 6.72% —> change MFC SF6 to 7% (it was 8%)
  • Probably a leak is present: to check when we are back. We can use the new R134a sniffer —> LPh+Fabio/Federico

Other

  • Change He GC bottle

 

ATLAS #2427

MDT

  • New pressure sensors with problem: need to change or reset when we are back. To prepare a list of sensors that stop to work or drift in order to have a clear overview —> Kacper
  • Presentation by Gianluca and Kacper on overshoot pressure problem:
    • ~ 30 mbar of overshoot and 15 hours to stabilise
    • overshoot is mitigated with higher Kc: doubling the Kc reduce of half the overshot
    • overshoot vs duration power-cut in the past —> Kacper
    • Gianluca checked and the PID are the same of last year
    • Roberto proposes also to lower the Ti to ~1
    • Roberto suggests that maybe the PID box changed —> Michal to investigate
    • Beatrice asked if there could be any correlation in the change of pneumatic pipes and FESTO Cabinet. To check speed of opening of regulation valve —> Michal
    • Tests to do next week: discussion offline to discuss the procedure —> Kacper, Gianluca, Beatrice, Roberto

TGC

  • To send an email to Leatitia for the safety form for preparing the intervention for when we are back —> Kacper/Gianluca

 

 

CMS #2426

RPC

  • Andrea increased flow in racks 84 and 86, probably problem in flowcell reading in UX. To verify when we are back —> Michal/Mara
  • Rack 80 ch 7 problem in reading: to check cables in flowcell when there is an access —> Andrea

From previous meeting:

  • with the power-cut some racks went underpressure -2 mbar, this is probably due to leaks in the detectors —> Andrea to prepare list of racks with this problem

 

LHCb #2430

RICH2

  • change pump membrane (pre-alarm of membrane pressure received, very good!)
  • New compressor module: Patrick to check with LHCb if it is possible to install in flammable room.

MWPC

  • change pump membrane

RICH1

  • to restart the system this week or beginning of next in order to try to understand if we take Air —> Patrick
  • Patrick proposed two tests:
    • Check Air intake with system running continuously
    • Check Air intake with stop and start of the system

General

  • To remember that there could be the possibility of having a RPC-like mixer to install in LHCb (CODEX-b project)

 

GENERAL

  • Presentation of 3D drawings by Leonardo:
    • 3 modules ready: mixer, exhaust and pump
    • now Leonardo is doing the purifier
    • to check mixer module and to send comment to Leonardo before next meeting —> everybody
    • missing list of material for mixer and exhaust —> Leonardo to understand if to do now or when we are back with help of technicians
    • LPh will send to Leonardo list of material for Exhaust module
  • Presentation of PROFIBUS by Patrick:
    • Two possible types of check:
      • with profibus trace: we check the voltages. They have to be ~ 5V
      • with multimeter: we check the resistances. They have to be < 20 Ohm.
      • the second test looks more easy for looking at possible problems
      • for both tests the gas system has to be OFF
      • good example of ALICE TRD in slides #6 and #7
      • for doing these type of test we need 2 people and 1 day per system
    • The profibus signal of ALICE TRD was very bad with old profibus module
    • With new profibus module the signal improved a lot but still some reflections visible
    • After the change of 9 connectors the signal is now good
    • A profibus connector costs ~30-40 chf (the one that are old and that could have problem)
    • We have 29 connectors for each system
    • To do a proposal for WinCC-OA errors from profibus —> Michal
    • For Profibus signal we can have maximum 32 stations without repeater: enough for LHC but not for CLOUD —> to investigate
    • Profibus check has to be done before and after the change of PLC but it is not enough
    • Need to test all resistance in profibus connectors. Not done for all system where we changed the PLC —> to plan to do
    • Test on Profibus status in OT in order to have a procedure of all tests needed
  • PGS list of sensors not working: missing ALICE
  • Two lists of warnings: one for the piquet (very important for operation and to anticipate interventi to avoid a stop of the system) and one for standard operation: Michal said it is possible —> to define what we want
    •  

From previous meeting:

  • Level of PGS sometimes is > 100% —> Michal to check

 

ACTIONS:

  • To check mixer module and to send comment to Leonardo before next meeting —> everybody
  • Kacper to prepare a table of ATLAS MDT pressure sensors with problem and keep it updated
  • To prepare a list of sensors not working or missing in PGS —> LPh
  • To send an email to Leatitia for the safety form for preparing the intervention for when we are back —> Kacper/Gianluca

From previous meeting:

  • To prepare a list of sensors not working or missing in PGS —> LPh
  • With the power-cut some racks went underpressure -2 mbar, this is probably due to leaks in the detectors —> Andrea to prepare list of racks with this problem
  • Excel file for alarms prepared weekly by Mara —> Experiment technical responsible to fill
There are minutes attached to this event. Show them.
    • 09:00 09:10
      SMS alarms from gas systems 10m
      Speaker: Mara Corbetta (Universite Claude Bernard Lyon I (FR))
    • 09:10 09:20
      Reminder of activities to do at the re-start 10m
      Speakers: Andrea D'Auria (CERN), Beatrice Mandelli (CERN), Frederic Merlet (CERN), Gianluca Rigoletti (Universite Claude Bernard Lyon I (FR)), Kacper Jerzy Kapusniak (CERN), Louis-Philippe De Menezes (CERN), Mara Corbetta (Universite Claude Bernard Lyon I (FR)), Michal Zbigniew Zimny (CERN), Patrick Carrie (CERN), Roberto Guida (CERN), Sabino Zagaria (Universita & INFN Pisa (IT))

      REMINDER for when we will be back

       

      ALICE

      CPV

      • We have 3 lines input that are ok, and 3 lines output: in one there is not the return. LPh contacted Yuri
      • electrical chassis of humidifier to finish
      • installation and commissioning of humidifier

      TPC

      • In the purifier the small thermic box triggered (also last year it happened) —> Patrick/LPh to check
      • Modification filter
      • New power supply and PLC (~week 5)

      TRD

      • Recuperation: missing two valves but there are present in the software
      • Validation of two reservoirs
      • Purifier: to review if keep the safety valve or the rupture disk
      • Purifier: safety valve at the input where we have Ar or Ar/H2 at high pressure —> LPh to check with safety

      MTR

      • Exhaust and humidifier modifications
      • Dummy chamber to complete installation

      TOF

      • Search for leak (~10-15 l/h)

       

       

      ATLAS

      MDT

      • Check overpressure in MDT chamber when there is a stop-start of the system
      • Only one convertisseur, one pump, etc from the system —> Patrick/Kacper to fix
      • Found a problem of grounding on pressure sensors: pressure sensors changed. Now to verify if the grounding problem is still there
      • Problem in the 2 vaisala of humidifier: several spikes present. Not clear why. Before the software upgrade they were not present

      RPC

      • Re-start of the gas system and related tests —> Kacper to check with Enrico
      • Fill 3rd purifier
      • IR analyser to move in exhaust module: to schedule later when there will be PLC change
      • dummy chamber pressure sensors + connectors to install
      • cabling/connectors for new racks

      TGC

      • Welding in nC5H12 room
      • To send an email to Leatitia for the safety form for preparing the intervention for when we are back —> Kacper/Gianluca

      MM

      • To start gas system mechanical and electrical
      • To provide MFCs ranges —> Beatrice

       

      CMS

      RPC

      • Finish to test regulation valves -> Beatrice
      • Michal changed the mother board in rack 67: now flow is 170 l/h (elog #2279) —> Andrea to check with new voegtlin red
      • Calibration of IR and GC for iC4H10 —> Mara
      • Change material of purifier 1 and 2 and fill 3rd (only when RPC will be in stop, probably September/October)
      • To continue dummy chamber installation (when there is access)
      • Andrea increased flow in racks 84 and 86, probably problem in flowcell reading in UX. To verify when we are back —> Michal/Mara
      • Rack 80 ch 7 problem in reading: to check cables in flowcell when there is an access —> Andrea
      • Missing one pressure sensor in rack xx and in rack 61 the pressure sensor show GEM pressure

      DT

      • Changed one motherboard in rack 61 but flow didn’t change (elog #2280) —> to try to recalibrate one set of flowcells
      • Difficult to re-start the gas system after power-cut —> to do some tests
      • Test for power cut and re-start for the exhaust module —> to do other tests

      GEM

      • Andrea prepared a drawing for ME0 and check for space in UX —> Andrea to send an email to Michele

      CSC

      • On-going regulation of PID for the exhaust —> Beatrice/Mara/Leonardo
      • Warning to set-up —> Beatrice/Michal/Roberto

      CF4 recovery

      • Install new storage battery —> Andrea
      • New MFC installation
      • New membrane installation
      • To fix leaks in column B of CF4 absorber
      • Test new compressor

      PGS

      • Check Freon supply > weight on side A/B (Mara)

      R134a

      • Restart tests

      GENERAL

      • Check of bande-chauffantes under floor and panel in

       

      LHCb

       

      MWPC

      • Problem in the GC analysis of the mixer: not visible the correct gas mixture with both LF and HF. Probably an hardware problem (connection, valve, etc) to find —> Mara
      • Leak in Mixer and Exhaust. Only in Exhaust found but not repaired. Other not yet founded. After finding the leaks, new weld to do —> Patrick

      RICH2

      • To inject Air —> Patrick to check with C. Frei
      • To add line for analysis “supply to detector”
      • GC analysis of recuperated CF4
      • All new upgrades of the system: IR, compressor, membrane
      • New PLC (week 1)

      SciFi

      • Problem on 3-ways valve —> Patrick

      PGS

      • Interlock of Ar/H2 does not work —> Michal/Patrick to check

       

      GENERAL

       

      • Alarms/Errors in power supplies for several gas systems —> Patrick
      • Profibus check has to be done before and after the change of PLC —> Michal
      • Check of all Profibus connectors —> Patrick
      • Aging test of: CF4 membrane, LHCb CO2 IR,  LHCb RI2 Coriolis, RPC new exhaust flowmeter, black flowmeters Omron (ordered by priority)
      • MFC calibration
      • Installation of pressure sensors for filters in exhaust and pump

       

       

    • 09:20 09:30
      M&O and project organization 10m
      Speakers: Beatrice Mandelli (CERN), Roberto Guida (CERN)
    • 09:30 09:40
      Overview of planning LS2 10m
      Speaker: Beatrice Mandelli (CERN)
    • 09:40 09:45
      Piquet interventions 5m

      http://eloggas.cern.ch/GAS_PIQUET/

      Speakers: Andrea D'Auria (CERN), Beatrice Mandelli (CERN), Frederic Merlet (CERN), Louis-Philippe De Menezes (CERN), Patrick Carrie (CERN), Kacper Jerzy Kapusniak (CERN), Michal Zbigniew Zimny (CERN)
    • 09:45 09:50
      Activities on-going 5m
      Speakers: Beatrice Mandelli (CERN), Frederic Merlet (CERN)
    • 09:50 10:10
      Plan of LS2 activities 20m
      Speakers: Beatrice Mandelli (CERN), Frederic Merlet (CERN), Roberto Guida (CERN)
    • 10:10 11:35
      Gas systems for the LHC experiments
      • 10:25
        ALICE 5m
        Speaker: Louis-Philippe De Menezes (CERN)
      • 10:30
        ATLAS 5m
        Speaker: Kacper Jerzy Kapusniak (CERN)
      • 10:35
        CMS 5m
        Speaker: Andrea D'Auria (CERN)
      • 10:40
        LHCb 5m
        Speaker: Patrick Carrie (CERN)
      • 10:45
        Gas analysis: standard H2O/O2 5m
        Speaker: Louis-Philippe De Menezes (CERN)
      • 10:50
        Gas web-page 5m

        http://test-gasmonitoring.web.cern.ch
        https://ep-dt-gasanalysis.web.cern.ch

        Speaker: Beatrice Mandelli (CERN)
      • 10:55
        Software 10m
        Speaker: Michal Zbigniew Zimny (CERN)
    • 11:35 12:25
      Gas System developments
      • 11:35
        Gas System Test-Bench 10m
        Speakers: Fabio Bordon (CERN), Sabino Zagaria (Universita & INFN Pisa (IT))
      • 11:45
        CMS CF4 recovery 10m
        Speakers: Fabio Bordon (CERN), Mara Corbetta (Universite Claude Bernard Lyon I (FR))
      • 11:55
        LHCb RICH2 CF4 recuperation 10m
        Speakers: Federico Consolaro (CERN), Mara Corbetta (Universite Claude Bernard Lyon I (FR))
      • 12:05
        R134a Recuperation 10m
        Speakers: Federico Consolaro (CERN), Gianluca Rigoletti (Universite Claude Bernard Lyon I (FR)), Ian Crotty (University of Wisconsin Madison (US)), Mara Corbetta (Universite Claude Bernard Lyon I (FR))
      • 12:15
        AoB 10m
        Speakers: Fabio Bordon (CERN), Federico Consolaro (CERN), Gianluca Rigoletti (Universite Claude Bernard Lyon I (FR)), Mara Corbetta (Universite Claude Bernard Lyon I (FR)), Sabino Zagaria (Universita & INFN Pisa (IT))
    • 12:25 12:35
      FSU & Jobs status 10m
      Speaker: Frederic Merlet (CERN)
    • 12:35 12:45
      AoB 10m
      Speakers: Andrea D'Auria (CERN), Beatrice Mandelli (CERN), Frederic Merlet (CERN), Louis-Philippe De Menezes (CERN), Patrick Carrie (CERN), Roberto Guida (CERN)